OpenStudy (anonymous):

I dont have alot of time so can you please help on this question! Find the surface are of a cylinder with a height of 10 in and a 2o in radius! Please help fast!

OpenStudy (unklerhaukus):

A cylinder has three surfaces; two circles, and a rectangle

OpenStudy (unklerhaukus):

If the radius of the circle is 20 find the area of the circles with the formula \[A_\text{circle}=\pi r^2\]

OpenStudy (anonymous):

I tried that and it gave me an answer that wasnt a choice, I don't know what im doing wrong :(

OpenStudy (unklerhaukus):

What area do you get for a circle ?

OpenStudy (unklerhaukus):

\[\pi\times(20)^2=\]